OpenStudy (anonymous):

What is the value of the y variable in the solution to the following system of equations? 3x - 6y = 12 -2x + 3y = 6

OpenStudy (phi):

simplify the first equation by dividing each term by 3 then multiply each term by +2 add equations solve for y

OpenStudy (campbell_st):

3x - 6y = 12 (1) -2x + 3x = 6 (2) multiply equation (2) by 3 and equation (1) by 2 and add to eliminate x 6x - 12y = 24 -6x + 9y = 18 -------------- then solve for y

OpenStudy (anonymous):

y = 2 ?

OpenStudy (campbell_st):

well adding the equations gives -3y = 42 solve for y

OpenStudy (anonymous):

y = 14